OpenStudy (anonymous):

Please help! Mary has 2/3 of a batch of cookies. She gives her mother 1/6 of her cookies. What is the fraction of the whole that her mother received?

OpenStudy (anonymous):

The question is basically asking "What is 1/6th of 2/3rds?" We can find this by multiplying the two fractions: \[\frac{1}{6} \times \frac{2}{3} = \frac{1 \times 2}{6 \times 3} = \frac{2}{18}=\frac{1}{9}\] Mary's mom has 1/9th of the total amount of cookies.

OpenStudy (anonymous):

Thank you so very much. Now it seems so simple....duh.